OpenStudy (anonymous):

Which of the following pairs of fractions are equivalent? Select ALL that apply A. 9/32 AND 18/64 B. 18/27 AND 24/36 C. 8/12 AND 6/10 D. 10/12 AND 12/15 E. 40/16 AND 25/10

OpenStudy (lucaz):

simplify the fractions

OpenStudy (lucaz):

\[\Large{\frac{18}{64}=\frac{9(2)}{32(2)}=\frac{9}{32}}\]

OpenStudy (anonymous):

thanks

OpenStudy (lucaz):

no problem